# Vascular Ageing and Smoking Habit Prediction via a Low-Cost Single-Lead ECG Module

### v2

@@ IV-B Data Pre-Processing
Segmentation: Each of the 42 examples/recordings were further divided into a number of five-second long (overlapping) segments with a stride of one second. This led to an increase in number of examples from 42 to 6131. Table I outlines the distribution of our dataset after segmentation.
Table I: Distribution of our dataset (after segmentation)
Class Label No. of Examples
Male $3455$
Female $2676$
Total $6131$
Male Smokers $2111$
Female Smokers $734$
Total Smokers $2845$
Male Non-Smokers $1344$
Female Non-Smokers $1942$
Total Non-Smokers $3286$
IV-C Feature Extraction
@@
We extracted a number of features, e.g., RR interval, QT interval, P wave duration, PP interval, TP interval, PR interval, T wave duration, ST segment, QRS complex duration, PR segment, etc., from the segmented ECG data (using the Neurokit2 library). The extraction process involved applying a discrete wavelet transform to the ECG signals, which helped focus on the various peaks in the PQRST wave (i.e., one cardiac cycle) and the corresponding intervals.
@@
Recall that each ECG segment is of duration 5 seconds, thus, each segment typically consists of about 5-10 cardiac cycles. Therefore, though we initially did feature extraction on a cardiac-cycle basis, we then averaged all the 13 extracted features across a segment in order to achieve reliable and stable features. We further observe that the values of the extracted features lie within the nominal range specified by the American heart association for the healthy young people [34]. Finally, Table II presents the values of the extracted features for one example from our dataset (i.e., corresponding to one segment of a male subject), for illustration purposes.
ECG Features Value (ms)
RR Interval $712.32$
RMSSD $65.59$
SDNN $52.40$
QT Interval $362.33$
P-wave Duration $101.47$
PP Interval $777.88$
TP Interval $211.74$
PR Interval $135.79$
T-wave Duration $176.13$
ST Segment $59.36$
QRS Complex Duration $126.84$
Corrected QT Interval $429.31$
PR Segment $34.32$
Table II: An illustration of the feature extraction process (i.e., one example from our dataset, corresponding to one segment of a male subject).
@@ V Machine, Deep and Transfer Learning Models Implemented
Machine learning models: For the vascular ageing prediction problem, we implemented the following ML models: linear regression, ridge regression (linear regression with regularization), decision tree regression, and random forest regression (an ensemble of decision trees). For smoking prediction problem, we implemented the following ML models: decision tree, random forest, Gaussian naive bayes, and XGBoost.
@@
Deep learning models: For the vascular ageing prediction problem, we implemented a 1D CNN (using the Keras library). The model utilizes 32 filters and a kernel size of 3, followed by a ReLU activation function. In the fully connected layers, a ”Dense” layer with 128 neurons and a ReLU activation are employed, followed by a dropout layer with a rate of 0.2 to mitigate overfitting. Lastly, in the output layer, a ”Dense” layer with a single neuron and a linear activation function is used for regression. The model is compiled using the Adam optimizer, with 220 epochs and a batch size of 23. For smoking prediction problem, we implemented a multi-layer perceptron (MLP) with one hidden layer having 10 neurons. We used the tanh activation function, and set the regularization parameter Alpha to 0.01. (We did feature selection and passed 8 most significant features to the MLP).
@@
Transfer learning model: For the vascular ageing prediction problem, we did transfer learning by utilizing ResNet18, a popular CNN model in computer vision community. The model has 18 layers, and takes as input the RGB images of size 224x224 pixels. The structure of ResNet-18 is as follows: i) convolutional layer (7x7, stride 2), for feature extraction; ii) max-pooling (3x3, stride 2), to reduce the spatial dimensions of the feature maps; iii) residual blocks (5 groups with 2 or 3 basic blocks each) with skip connections [^2] Each BasicBlock layer comprises a convolutional layer, followed by batch normalization, followed by a ReLU activation function.; iv) global average pooling, to further reduce the spatial dimensions of the feature maps; v) a fully connected layer with 512 input features, followed by a ReLU activation, an output layer with a single neuron to predict the vascular age of the input sample.
@@
Training of ML, DL and TL models: The dataset (consisting of 6131 samples) was divided into three parts. 60% (i.e., 3679 samples) of the dataset was used for training, 20% (i.e., 1226 samples) for validation, and remaining 20% (i.e., 1226 samples) for testing. For each method, the best combination of hyperparameters was identified by conducting a grid search. For transfer learning, we translated the 1D ECG segments into spectrograms and fed them to the ResNet18 model.
@@ VI-A Performance Metrics
For the regression problem (of vascular aging prediction), we utilize the mean squared error (MSE) and $R^{2}$ score as performance metrics. Note that a higher $R^{2}$ score means a better fit and a lower MSE means a better model predictive accuracy. For the classification problem (of smoking habit prediction), we construct confusion matrices which provide details such as true positives, true negatives, false positives, and false negatives.
@@ VI-B Vascular Ageing Prediction
We begin by doing the feature correlation analysis. Fig. 5 reveals that vascular ageing is agnostic to the following features: BMI, sleep duration, RMSSD, SDNN and PR interval. Similarly, we learn from Fig. 5 that following features are most relevant to the vascular ageing: heart rate, P-wave duration, QT interval, PP interval, and T-wave duration.
@@
Figure 5: Feature correlation analysis (target variable is age).
Among all the ML models, the linear models, i.e., linear regression model and its variant ridge regression performed the worst with an MSE of 5.79, $R^{2}$ score of 0.337, and MSE of 5.24, $R^{2}$ score of 0.37, respectively. This hints at a possible non-linear relationship between the ECG-based features and the vascular age. This hypothesis is later verified as the decision tree model and random forest model (an ensemble of decision trees) performed very well with a very small MSE of 0.258, high $R^{2}$ score of 0.97, and small MSE of 0.07, high $R^{2}$ score of 0.99, respectively. On the deep learning front, the 1D CNN also performed well with an MSE of 0.251 and an $R^{2}$ score of 0.971 (at a cost of more computational complexity). Finally, ResNet18 model achieved an MSE of 1.191, and a $R^{2}$ score of 0.86. Table III (bottom half) summarizes the MSE and $R^{2}$ results for all 6 of our models.
Table III also highlights that our ML models (random forests and decision tree) and DL model (1D CNN) provide superior MSE and $R^{2}$ performance compared to the 12-lead ECG-based state-of-the-art methods for vascular aging prediction.
@@
Starc et al. [35]
2012
12-lead ECG
Multiple Linear Regression
3-100
R2=0.76
Lima et al. [16]
2021
12-lead ECG
Residual Network
20-80
MAE=8.83
Chang et al. [18]
2022
12 Lead ECG
Deep Learning
20-80
MAE=6.899 R2=0.822
Zvuloni et al. [36]
2023
12-lead ECG
FE and DL
20-85
MAEFE=10.6 R2FE=0.60 MAEDL=6.32 R2DL=0.83
Ours 2023 1-lead ECG Random Forest 18-30 MSE=0.07; R2=0.99
Ours
2023
1-lead ECG
Decision Tree
18-30
MSE=0.258; R2=0.97
Ours
2023
1-lead ECG
Ridge Regression
18-30
MSE=5.24; R2=0.37
Ours
2023
1-lead ECG
Linear Regression
18-30
MSE=5.79; R2=0.337
Ours
2023
1-lead ECG
1D CNN
18-30
MSE=0.251; R2=0.971
Ours
2023
1-lead ECG
ResNet18
18-30
MSE=1.191; R2=0.86
Table III: MSE and $R^{2}$ accuracy of our ML, DL and TL models and their comparison with the previous work.
Fig. 6 plots the histograms of the error in vascular aging prediction for all 6 models we have implemented. We observe that the random forest, the decision tree, and 1D CNN have the smallest error spread among all, as expected.
@@ 12 Lead ECG
VI-C Smoking Habits Prediction
We again do feature correlation analysis first. Fig. 7 reveals that smoking makes the least impact on the following features: QRS complex duration, PR interval, QTc, and heart disease history in the family. Similarly, we learn from Fig. 7 that following features alter the most due to smoking: gender, ST segment, RR interval, systolic blood pressure, and heart rate.
@@ VI-C Smoking Habits Prediction
Figure 7: Feature correlation analysis (target variable: smoking).
Now, the first problem at hand here is the binary classification problem where we aim to differentiate between a smoker and non-smoker using the ECG-based features and other features. For this problem, XGBoost stood out among all four models with an accuracy of 96.5%. The other three models, i.e., Gaussian naive bayes, Decision tree, Random forest achieved relatively low accuracy of 66.2%, 67.4%, 75.2%, respectively. Table IV provides the detailed confusion matrices for all the four classifiers, for this problem.
The next problem at hand is the 4-class classification problem that aims to differentiate between male smokers, female smokers, male non-smokers, and female non-smokers, using the ECG-based features and other features. Among all four models, the multi-layer perceptron performed the best with an accuracy of 97.5%. The decision tree was next in line with an accuracy of 90.5%. The remaining two models, i.e., Gaussian naive bayes, Random forest achieved relatively low accuracy of 66.0%, 87.3%, respectively. Table V provides the detailed confusion matrices for all the four classifiers, for this problem.
Algorithm True Predicted
Non-Smoker
Smoker
Gaussian Naive Bayes
Non-Smoker
Smoker
526
372
42 287
Decision Tree
Non-Smoker
Smoker
363
194
205 465
Random Forest
Non-Smoker
Smoker
374
110
194 549
XGBoost
Non-Smoker
Smoker
541
15
27 644
Table IV: Confusion matrices of the four ML classifiers for the binary classification problem (that aims to differentiate between a smoker and a non-smoker).
Algorithm True Predicted
Male Smoker
Female Smoker
Male Non-Smoker
Female Non-Smoker
Gaussian Naive Bayes
Male Smoker
Female Smoker
Male Non-Smoker
Female Non-Smoker
345
89
2
65
42 128
0
67
39 41 130
47
6 15 4 207
Decision Tree
Male Smoker
Female Smoker
Male Non-Smoker
Female Non-Smoker
832
66
0
75
0 467
0
55
15 5 290
6
6 4 1 631
Random Forest
Male Smoker
Female Smoker
Male Non-Smoker
Female Non-Smoker
407
98
4
0
0 160
0
4
0 0 123
0
25 15 9 382
Multi Layer Perceptron
Male Smoker
Female Smoker
Male Non-Smoker
Female Non-Smoker
532
10
9
2
8 832
12
0
2 11 742
1
0 0 4 288
Table V: Confusion matrices of the four classifiers for the 4-class classification problem (that aims to differentiate between male smoker, female smoker, male non-smoker, and female non-smoker).
VII Conclusion
This paper presented a novel method to predict the vascular age and smoking status of a person using a custom-built low-cost single-lead ECG sensor module. We constructed our own dataset by collecting ECG data and metadata from 42 apparently healthy subjects (male, female, smoker, non-smoker). We passed the pre-processed and segmented dataset to a number of ML models, a 1D CNN, an MLP, and ResNet18 model. For vascular ageing prediction problem, Random Forest method performed the best with an $R^{2}$ score of 0.99, and MSE of 0.07. For the 2-class smoking classification problem, XGBoost method stood out with an accuracy of 96.5%. Finally, for the 4-class smoking classification problem, MLP method achieved the best accuracy of 97.5%.
This work is aligned with the sustainable development goals (3 and 10) of the United Nations which aim to provide low-cost but quality healthcare sensing solutions to the unprivileged (e.g., people in developing countries and people in remote areas). Furthermore, this work opens the door for innovation in wearable devices (e.g., smartwatches and wristbands), enabling accurate and real-time prediction of the vascular age using a single-lead (lead-I) ECG sensor.
As for the future work, we note that our dataset consists of ECG data from apparently healthy young subjects only. Therefore, further efforts are required to collect large amounts of data from a more diverse population (with a much wider age range and with a wide range of cardiovascular diseases). Such dataset once collected, could also help understanding the implications of smoking on cardiovascular health. Further, there is a dire need to validate our results by comparing them with the established classical score-based methods. This investigation will provide valuable insights about the generalization capability of our models in real-world healthcare settings.

### v3

@@ IV-B Data Pre-Processing
IV-C Creation of 2nd Dataset via Segmentation
Inline with the established practice of data augmentation through segmentation in the field of supervised learning, each of the 42 examples/recordings were further divided into a number of five-second long (overlapping) segments with a stride of one second. This led to another labeled dataset with 6,131 data points. Table II outlines the distribution of the two custom datasets, i.e., the original/unsegmented dataset, and segmented dataset.
Unsegmented Segmented
Male 26 $3455$
Female 16 $2676$
Total 42 $6131$
Male Smokers 16 $2111$
Female Smokers 4 $734$
Total Smokers 20 $2845$
Male Non-Smokers 10 $1344$
Female Non-Smokers 12 $1942$
Total Non-Smokers 22 $3286$
Table II: Key statistics of the two custom datasets.
IV-D Feature Extraction
@@
We extracted a number of features, e.g., RR interval, QT interval, P wave duration, PP interval, TP interval, PR interval, T wave duration, ST segment, QRS complex duration, PR segment, etc., from both segmented ECG dataset and unsegmented ECG dataset (using the Neurokit2 library). The extraction process involved applying a discrete wavelet transform to the ECG signals, which helped focus on the various peaks in the PQRST wave (i.e., one cardiac cycle) and the corresponding intervals.
@@
We make the following observation regarding the segmented dataset. Recall that each ECG segment is of duration 5 seconds, thus, each segment typically consists of about 5-10 cardiac cycles. Therefore, though we initially did feature extraction on a cardiac-cycle basis, we then averaged all the 13 extracted features across a segment in order to achieve reliable and stable features.
Table III presents the values of the extracted features for one example from each of the two custom datasets (i.e., segmented and unsegmented), for illustration purposes. We observe that the values of the extracted features lie within the nominal range specified by the American heart association for the healthy young people [42].
ECG Features Segmented Unsegmented
RR Interval $712.32$ 701.15
RMSSD $65.59$ 26.78
SDNN $52.40$ 26.12
QT Interval $362.33$ 281.24
P-wave Duration $101.47$ 75.7
PP Interval $777.88$ 657.02
PT Interval $211.74$ 421.69
PR Interval $135.79$ 230.45
T-wave Duration $176.13$ 129.21
ST Segment $59.36$ 204.78
QRS Complex Duration $126.84$ 76.48
Corrected QT Interval $429.31$ 410.14
PR Segment $34.32$ 30.29
Table III: An illustration of the feature extraction process (i.e., one example from each of the two datasets, i.e., segmented and unsegmented datasets). Values are in milli sec.
@@ V Machine, Deep and Transfer Learning Models Implemented
Keeping in mind the very small size of our custom ECG dataset (consisting of 42 subjects only), the high-complexity AI models such as transformers, deep neural networks, large language models are ruled out. This leaves us with classical ML models which typically require a few hundred of examples in order to learn the regression problem at hand.
V-A Machine and deep learning models
@@
Machine learning models: For the vascular ageing prediction problem, we implemented the following ML models: linear regression, ridge regression (linear regression with regularization), decision tree regression, and random forest regression (an ensemble of decision trees).
@@
Deep learning models: For the vascular ageing prediction problem, we implemented a 1D CNN (using the Keras library). The model utilizes 32 filters and a kernel size of 3, followed by a ReLU activation function. In the fully connected layers, a dense layer with 128 neurons and a ReLU activation are employed, followed by a dropout layer with a rate of 0.2 to mitigate overfitting. Lastly, in the output layer, a dense layer with a single neuron and a linear activation function is used for regression. The model is compiled using the Adam optimizer, with 220 epochs and a batch size of 23.
V-B Transfer learning
For the vascular ageing prediction problem, we implemented two transfer learning models, a ResNet18 model for segmented ECG dataset, and a random forest model for unsegmented ECG dataset.
@@ V-B Transfer learning
First, we passed the segmented ECG dataset through ResNet18, a popular CNN model in computer vision community, in order to do transfer learning. The model has 18 layers, and takes as input the RGB images of size 224x224 pixels. The structure of ResNet-18 is as follows: i) convolutional layer (7x7, stride 2), for feature extraction; ii) max-pooling (3x3, stride 2), to reduce the spatial dimensions of the feature maps; iii) residual blocks (5 groups with 2 or 3 basic blocks each) with skip connections [^2] Each BasicBlock layer comprises a convolutional layer, followed by batch normalization, followed by a ReLU activation function.; iv) global average pooling, to further reduce the spatial dimensions of the feature maps; v) a fully connected layer with 512 input features, followed by a ReLU activation, an output layer with a single neuron to predict the vascular age of the input sample.
Secondly, we trained our custom-built AI models (i.e., linear regression, ridge regression, decision tree, random forest and 1D-CNN) on a photoplethysmography (PPG)-based vascular age estimation dataset of [34]. Later, we fine-tuned and tested this random forest model on our unsegmented ECG dataset that was originally pre-trained on PPG data. This approach has many different names in literature, e.g., domain adaptation, transfer learning, representation learning, meta learning, etc.
V-C Training of ML, DL and TL models
@@
The segmented ECG dataset (consisting of 6131 samples) was divided into three parts. 60% (i.e., 3679 samples) of the dataset was used for training, 20% (i.e., 1226 samples) for validation, and remaining 20% (i.e., 1226 samples) for testing. For each method, the best combination of hyperparameters was identified by conducting a grid search. For transfer learning, we translated the 1D ECG segments into spectrograms and fed them to the ResNet18 model. On the other hand, for the unsegmented ECG dataset, we do $k$-fold cross-validation with $k=5$.
@@ VI-A Performance Metrics
For the regression problem (of vascular aging prediction), we utilize the mean squared error (MSE) and $R^{2}$ score as performance metrics. Note that a higher $R^{2}$ score means a better fit and a lower MSE means a better model predictive accuracy.
@@ VI-B Vascular Ageing Prediction
Explainable AI analysis: We begin by doing the feature correlation analysis in spirit of explainable AI (xAI), for one example from each of the two datasets, i.e., segmented ECG dataset and unsegmented ECG dataset. We learn from Fig. 4 that feature extraction from both segmented and unsegmented ECG datasets yields very similar results. That is, both Fig. 4 (a) and Fig. 4 (b) illustrate that the following ECG features are most relevant to the vascular ageing: heart rate, P-wave duration, QT interval, PP interval, and T-wave duration. Similarly, both Fig. 4 (a) and Fig. 4 (b) reveal that vascular ageing is agnostic to the following ECG and demographic features: BMI, sleep duration, RMSSD, and SDNN.
However, we observe that despite lots of agreement between Fig. 4 (a) and Fig. 4 (b), there are some differences as well. For example, ST segment is considered as an important ECG feature by the feature extraction process done on unsegmented dataset, but not by segmented dataset. As another example, PR segment is considered as an important ECG feature by the feature extraction process done on segmented dataset, but not by unsegmented. We believe this minor discrepancy is an artifact due to small size of our custom dataset. But the overall agreement between Fig. 4 (a) and Fig. 4 (b) points to the legitimacy of the segmentation process.
(a) Using segmented ECG dataset.
(b) Using unsegmented ECG dataset.
@@
Figure 4: Explainable AI based feature correlation analysis. (target variable is age).
AI model Segmented (S) Unsegmented (US) Unsegmented+transfer learning (US+TL)
Linear regression MSE=5.79; $R^{2}$=0.337 MSE = 14.13; $R^{2}$ = -1.92 MSE= 4.36; $R^{2}$= 0.45
Ridge regression MSE=5.24; $R^{2}$=0.37 MSE = 18.19; $R^{2}$ = -2.76 MSE = 4.39; $R^{2}$ = 0.45
Decision tree MSE=0.258; $R^{2}$=0.97 MSE = 6.13; $R^{2}$ = -0.26 MSE= 4.28; $R^{2}$=0.459
Random forest MSE=0.07; $R^{2}$=0.99 MSE = 3.56; $R^{2}$ = 0.26 MSE = 0.99; $R^{2}$= 0.87
1D-CNN MSE=0.251; $R^{2}$=0.971 MSE = 4.24; $R^{2}$ = 0.12 MSE= 28.2; $R^{2}$= -2.564
Table IV: Performance of our AI models for vascular ageing prediction, across the three learning scenarios.
Table IV provides a detailed performance comparison of all the AI models across the three distinct learning scenarios that we have implemented for vascular age prediction. Specifically, we have evaluated our AI models: 1) on the segmented ECG dataset, 2) on the unsegmented (original) ECG dataset as is, and 3) for the scenario where our models are first pre-trained on the PPG dataset of [34], and later fine-tuned and tested on our custom unsegmented ECG dataset. For the first learning scenario where we evaluate all the models on segmented ECG dataset, the linear ML models, i.e., linear regression model and its variant ridge regression perform poorly which hints at a possible non-linear relationship between the ECG-based features and the vascular age. This hypothesis is later verified as the decision tree model and random forest model (an ensemble of decision trees) perform very well with a smaller MSE and higher $R^{2}$ score. On the deep learning front, the 1D CNN also performs well but at a cost of more computational complexity. Next, for the second learning scenario where we evaluate all the models on unsegmented ECG dataset, we notice that the linear regression and ridge regression collapse because of the small size of the unsegmented dataset and because of non-linear relationship between the ECG-based features and the vascular age. Finally, for the third scenario where we fine-tune and test all the models (that were pre-trained on PPG dataset of [34]) on unsegmented ECG dataset, all models perform decently, but the 1D-CNN also collapses. This is perhaps due to the small size of our unsegmented ECG dataset that we use for fine-tuning purposes. All in all, the random forest model comes out as a winner across all three learning scenarios. That is, the random forest model achieves an MSE of 0.07 and $R^{2}$ of 0.99, MSE of 3.56 and $R^{2}$ of 0.26, MSE of 0.99 and $R^{2}$ of 0.87, for segmented ECG dataset, for unsegmented ECG dataset, and for transfer learning scenario, respectively. The superior performance of the random forest is because of the following: i) it is an ensemble method so it produces more reliable aggregate decisions due to so many decision tree at the backend, and ii) its complexity is well-matched with the size of our unsegmented ECG dataset.
VI-C Comparison with related works on vascular ageing prediction
Table V provides a detailed performance comparison of the top performing random forest model (for all three learning scenarios) against the related works on vascular ageing estimation. Table V highlights that our random forest model provides MSE and $R^{2}$ performance that is at par with the performance of 12-lead ECG-based state-of-the-art methods for vascular aging prediction.
@@ VI-C Comparison with related works on vascular ageing prediction
Starc et al. [43] 2012 12-lead ECG Multiple Linear Regression 3-100 R2=0.76
Lima et al. [16] 2021 12-lead ECG Residual Network 20-80 MAE=8.83
Chang et al. [18] 2022 12 Lead ECG Deep Learning 20-80 MAE=6.899 R2=0.822
Zvuloni et al. [44] 2023 12-lead ECG FE and DL 20-85 MAEFE=10.6 R2FE=0.60 MAEDL=6.32 R2DL=0.83
Ours 2024 1-lead ECG Random Forest (S) 18-30 MSE=0.07; R2=0.99
Ours 2024 1-lead ECG Random Forest (US) 18-30 MSE=3.56; R2=0.26
Ours 2024 1-lead ECG Random Forest (US+TL) 18-30 MSE=0.99; R2=0.87
Table V: MSE and $R^{2}$ accuracy of our ML, DL and TL models and their comparison with the previous work.
Fig. 5 plots the histograms of the error in vascular aging prediction for all models we have implemented, for both segmented and unsegmented ECG datasets. For the segmented ECG dataset, we observe that the random forest, the decision tree, and 1D CNN have the smallest error spread among all, as expected. For the unsegmented ECG dataset, the histograms don’t reveal much about the error spread and error trends, due to small size of the dataset.
(a) Using segmented ECG dataset.
(b) Using unsegmented ECG dataset.
@@
VI-D Effect of Light-but-Habitual Smoking on ECG
Explainable AI analysis: We again do feature correlation analysis in spirit of xAI, by taking one example from each of the two datasets, i.e., segmented and unsegmented datasets. We learn from Fig. 6 that feature extraction from both datasets, i.e., segmented ECG dataset and unsegmented ECG dataset, leads to findings that are mostly in agreement. That is, the following ECG and demographic features show the highest correlation with smoking: heart rate, P wave duration, systolic blood pressure, BMI, for both datasets. Further, the following ECG and demographic features show the highest negative correlation with smoking: gender, ST segment, RR interval, PP interval, PT (or TP) interval, for both datasets. However, we note that there is a slight disagreement between the two datasets regarding a few features, e.g., regarding QRS complex duration whereby the unsegmented dataset considers it important, while segmented dataset doesn’t. Similarly, sleep duration is important for segmented dataset but not for unsegmented dataset. Furthermore, QT interval is deemed more important by unsegmented dataset, while it is considered less important by segmented dataset. We believe this discrepancy is due to small size of our custom dataset. Overall, a good agreement between Fig. 6 (a) and Fig. 6 (b) validates the segmentation process carried out in this work.
(a) Using segmented ECG dataset.
(b) Using unsegmented ECG dataset.
@@ VI-D Effect of Light-but-Habitual Smoking on ECG
Figure 6: Explainable AI based feature correlation analysis. (target variable is smoking).
Next, to analyze the effect of smoking on ECG features at an aggregate level, we did feature extraction from the full unsegmented dataset. Fig. 7 plots the mean values and standard deviations of 11 most relevant ECG features for both smoker and non-smoker groups. We notice that the following features show noticeable changes (in terms of mean and standard deviation) across the two classes (i.e., smokers and non-smokers): RMSSD, and SDNN. This is followed by the following features: sleep duration and BMI. The following ECG features register mediocre changes across the two classes: systolic blood pressure, heart rate, QT interval, PP interval, and ST segment. On the other hand, the following ECG features remain unchanged across the two classes: diastolic blood pressure, and PT interval. Thus, even habitual-but-light smoking causes changes/abnormalities in ECG traces, as shown by the changes in the feature values in Fig. 7. This could in term help do early prediction of the risk of cardiovascular diseases among the smokers group.
Figure 7: Impact of smoking on ECG features.
VII Data availability statement
The custom ECG dataset used and analysed during this research study is available from the corresponding author on reasonable request.
VIII Declarations
VIII-A Funding and/or Conflicts of Interests/Competing Interests
Funding: This work was supported in part by the UK Engineering and Physical Sciences Research Council (EPSRC) grants: EP/X040518/1 and EP/T021020/1. Conflict of Interest: The authors declare that they have no conflict of interest.
IX Conclusion
This paper predicted the vascular age of a person and studied smoking-induced changes in ECG using a custom-built low-cost single-lead ECG sensor module. We constructed our own custom dataset by collecting ECG data and metadata from 42 apparently healthy young subjects (both smokers and non-smokers). We passed the segmented and unsegmented datasets to a number of AI models. We also did transfer learning whereby we pre-trained our models on a public PPG dataset, and later, fine-tuned and evaluated them on our unsegmented ECG dataset. The random forest model outperformed all other models and previous works by achieving a mean squared error (MSE) of 0.07 and cofficient of determination $R^{2}$ of 0.99, MSE of 3.56 and $R^{2}$ of 0.26, MSE of 0.99 and $R^{2}$ of 0.87, for segmented ECG dataset, for unsegmented ECG dataset, and for transfer learning scenario, respectively. Finally, we utilized the explainable AI framework to identify the most relevant ECG features that change due to smoking.
This work is aligned with the sustainable development goals (3 and 10) of the United Nations which aim to provide low-cost but quality healthcare sensing solutions to the unprivileged (e.g., people in developing countries and people in remote areas). Furthermore, this work opens the door for innovation in wearable devices (e.g., smartwatches and wristbands), enabling accurate and real-time prediction of the vascular age using a single-lead (lead-I) ECG sensor. In addition, this work also finds its applications in the broad domain of forensic science.
Finally, we note that our dataset consists of resting-state ECG data collected from apparently healthy young subjects only. Therefore, further efforts are required to collect large amounts of data from a more diverse population (e.g., with different ages, smoking habits, cardiovascular diseases, varying time gaps between last smoking instance and the ECG measurement etc.). Such dataset will help understand the more fine-grained implications of smoking on cardiovascular health, as well as provide valuable insights about the generalization capability of the AI models in real-world settings.